OpenStudy (zyberg):

How to solve without graphing equation: 1/sqrt{x} + 3sqrt{x} = 4? Roots are: 1/9 and 1.

OpenStudy (zyberg):

\(\frac{1}{\sqrt{x}} + 3\sqrt{x} = 4\)

OpenStudy (jango_in_dtown):

so 1+3x=4 root x now square both sides and you get (1+3x)^2=16x or, 9x^2+6x+1-16x=0 or, 9x^2-10x+1=0 or, 9x^2-9x-x+1=0 or, 9x(x-1)-1(x-1)=0 or, (x-1)(9x-1)=0 orm x=1, x=1/9

OpenStudy (jango_in_dtown):

@Zyberg